# Deep Pre-trained Time Series Features for Tree Species Classification in the Dutch Forest Inventory

### v1

@@ A.1.1 Results of models using only S1 and S2 data for fine-tuning
Table A.2: Classification performance comparison between MLP and RF classifiers with deep features (in %). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance between MLP and RF classifier per dataset type.
Results
Dataset type Data MLP RF
size classifier classifier
Overall Accuracy NFI data (7 classes) 1480 78.79 ± 1.01 75.47 ± 1.15
NFI data (13 classes) 1462 69.39 ± 1.65 67.77 ± 2.38
Francini data (7 classes) 13,790 98.08 ± 0.09 94.96 ± 0.58
F1 Score NFI data (7 classes) 1480 70.57 ± 2.20 63.83 ± 1.21
NFI data (13 classes) 1462 54.91 ± 4.29 48.24 ± 4.31
Francini data (7 classes) 13,790 98.07 ± 0.09 94.94 ± 0.58
Comparison between all bands and only S1 and S2 bands
Table A.3 compares models using all satellite data against those using only S1 and S2 data with MLP classifier. Interestingly, models with only S1 and S2 data achieved higher accuracy with NFI data, while no significant difference was observed with Francini data.
@@ Results
Table A.3: Classification performance comparison between using all available bands versus using only S1 and S2 bands (in %). Each model was run five times with mean and standard deviation (±) reported. Bold values indicate best performance between the two band configurations for each dataset type.
MLP classifier
Dataset type Data size All bands S1S2 only
Overall Accuracy NFI data (7 classes) 1480 77.13 ± 1.17 78.79 ± 1.01
NFI data (13 classes) 1462 67.99 ± 1.33 69.39 ± 1.65
Francini data (7 classes) 13,790 98.21 ± 0.20 98.08 ± 0.09
F1 Score NFI data (7 classes) 1480 68.24 ± 2.85 70.57 ± 2.20
NFI data (13 classes) 1462 52.45 ± 2.25 54.91 ± 4.29
Francini data (7 classes) 13,790 98.20 ± 0.20 98.07 ± 0.09

### v2

@@ A.3.1 Results of models using only S-1 and S-2 data for fine-tuning and embeddings extraction
Table A.3: Classification performance comparison between MLP and RF classifiers using fine-tuned Presto embeddings with only S-1 and S-2 inputs (in %). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance between the MLP and RF classifiers per dataset type.
Fine-tune Presto embeddings
Dataset type Data size MLP RF
Overall Accuracy SIMB (7 classes) 1,479 74.98 ± 1.78 74.01 ± 1.83
COMB (13 classes) 1,462 67.09 ± 0.60 65.42 ± 1.92
SIBA (7 classes) 13,790 92.23 ± 0.48 92.31 ± 0.18
F1 Macro SIMB (7 classes) 1,479 65.09 ± 2.97 60.00 ± 2.12
COMB (13 classes) 1,462 50.82 ± 3.87 44.67 ± 3.25
SIBA (7 classes) 13,790 91.91 ± 0.54 91.95 ± 0.24
F1 Weighted SIMB (7 classes) 1,479 74.49 ± 1.88 72.96 ± 1.87
COMB (13 classes) 1,462 65.00 ± 1.28 61.32 ± 2.10
SIBA (7 classes) 13,790 92.20 ± 0.49 92.23 ± 0.20
Comparison between all bands and only S-1 and S-2 bands
Table A.4 compares fine-tuned Presto embeddings obtained using all available satellite bands with those using only S-1 and S-2 data, evaluated with the MLP classifier. Overall, performance differences between the two band configurations are small across datasets, indicating that excluding auxiliary climatic (ERA5) and topographic (SRTM) information does not substantially degrade classification performance for this task.
For the SIMB dataset, using only S-1 and S-2 bands yields slightly higher accuracy and F1 scores, whereas for the COMB dataset both configurations perform comparably. In contrast, for the SIBA dataset, which is larger and more balanced, the use of all available bands results in marginally higher performance across all metrics. These results suggest that optical and radar data alone are sufficient to capture most discriminative information, while auxiliary bands provide limited additional benefit that appears to be dataset-dependent.
@@
Table A.4: Classification performance comparison between using all available bands and using only S-1 and S-2 bands for fine-tuned Presto embeddings with an MLP classifier (in %). Each model was run five times and the mean and standard deviation (±) are reported. Bold values indicate the best performance between the two band configurations for each dataset type.
MLP for fine-tune embeddings
Dataset type Data size All bands S-1 + S-2 only
Overall Accuracy SIMB (7 classes) 1,479 73.67 ± 1.29 74.98 ± 1.78
COMB (13 classes) 1,462 67.40 ± 1.47 67.09 ± 0.60
SIBA (7 classes) 13,790 93.54 ± 1.27 92.23 ± 0.48
F1 Macro SIMB (7 classes) 1,479 62.29 ± 4.60 65.09 ± 2.97
COMB (13 classes) 1,462 50.76 ± 2.23 50.82 ± 3.87
SIBA (7 classes) 13,790 93.54 ± 1.26 91.91 ± 0.54
F1 Weighted SIMB (7 classes) 1,479 73.05 ± 1.52 74.49 ± 1.88
COMB (13 classes) 1,462 65.51 ± 1.29 65.00 ± 1.28
SIBA (7 classes) 13,790 93.54 ± 1.26 92.20 ± 0.49
Confusion Matrix Comparison
Sets of confusion matrices for SIBA and COMB datasets show per-class accuracy improvements from deep embeddings, especially with fine-tuned Presto and the MLP classifier.
Figure 3 presents confusion matrices for the COMB dataset across six model configurations, combining different feature representations (fine-tuned deep embeddings, frozen deep embeddings, harmonic + medoid feature) and classifiers (MLP and Random Forest). This appendix provides a comparative description of class-level behavior across these configurations, complementing the detailed discussion of the fine-tuned MLP results in Section 5.3.
Across all models, broadleaved species exhibit substantially higher confusion than conifer species. In particular, Q. robur/petraea consistently acts as a dominant prediction class, absorbing misclassifications from several less frequent broadleaved species such as Alnus spp., Fraxinus spp., and Populus spp.. This pattern is visible irrespective of classifier choice, indicating that class imbalance and overlapping spectral–phenological characteristics dominate model behavior more strongly than the specific learning algorithm.
Comparing classifiers, RF models generally show more fragmented confusion patterns than the fine-tuned MLP. While RF-based models correctly identify dominant classes such as P. sylvestris and Q. robur/petraea, they exhibit reduced sensitivity to minority classes, with many rare species being almost entirely misclassified into majority classes. This effect is particularly evident for Alnus spp. and Fraxinus spp., which achieve very low recall across all RF configurations.
Feature representation also plays a key role. Fine-tuned deep embeddings yield the most diagonally concentrated confusion matrix, especially for conifer species. P. sylvestris and P. menziesii are consistently well separated under fine-tuned representations, whereas frozen deep features (Presto, AEF, and TESSERA) show increased confusion between conifer taxa, particularly between Pinus spp. and P. sylvestris. This suggests that task-specific fine-tuning is critical for capturing subtle intra-genus differences.
Models based on harmonic and medoid features display the weakest class separation overall. While they retain some ability to discriminate coarse functional groups (broadleaved vs. conifer), fine-grained discrimination among broadleaved species is limited.
Finally, large-scale pretrained representations (AEF and TESSERA) exhibit mixed behavior. While they perform competitively for dominant classes, they do not consistently improve minority-class recognition in the absence of fine-tuning. In several cases, their confusion patterns resemble those of frozen Presto embeddings, indicating that representation scale alone does not resolve class imbalance or inter-species similarity issues in the COMB dataset.
Overall, the confusion matrices for COMB reinforce three key observations: (i) class imbalance strongly shapes prediction outcomes across all models, (ii) fine-tuning deep embeddings improves intra-group species discrimination, and (iii) classifier choice has a secondary effect compared to data characteristics and feature representation. These findings support the conclusions drawn in the main text regarding the importance of data quality and class balance for operational NFI species mapping.
Figure 3: Confusion matrices for Complex & Imbalanced (COMB) datasets. Top left: Fine- tune Presto with MLP Classifier, Top right: Presto (Frozen Presto features), Middle left: Fine-tune Presto (Presto features), Middle right: Harmonic + medoid feature, Bottom left: AEF, Bottom right: TESSERA. Each cell has average count and recall percentage.
Figure 4 shows confusion matrices for the SIBA dataset across six model configurations, combining different feature representations and classifiers. In contrast to the COMB dataset, SIBA exhibits markedly stronger class separability across all models, reflecting its larger sample size and substantially improved class balance.
Across configurations, most classes achieve high diagonal dominance, indicating strong recall and precision for both broadleaved and conifer species groups. Fine-tuned Presto with an MLP classifier yields near-optimal performance for all classes, with precision values exceeding 90 % for most species and limited off-diagonal confusion. This confirms that, under favorable data conditions, fine-tuned deep embeddings can effectively capture inter-species differences in seasonal and structural signals.
RF classifiers show slightly reduced performance compared to the fine-tuned MLP but retain strong class-level behavior. In particular, dominant classes such as Beech, Larix, and Populus are consistently well identified across all feature representations. Residual confusion is primarily concentrated among broadleaved classes, especially between Other broadleaf and Quercus, suggesting remaining spectral and phenological overlap even in a balanced dataset.
Frozen deep feature representations (Presto, AEF, and TESSERA) display broadly similar confusion patterns. While they maintain high accuracy for majority classes, they exhibit increased confusion for classes with overlapping ecological traits. For instance, Dark conifer shows increased misclassification into Pinus and Larix under frozen representations, indicating that fine-tuning remains important for resolving subtle intra-group distinctions.
Models based on harmonic and medoid features show the weakest overall class separation, although their performance remains notably stronger than in the COMB dataset. This suggests that the improved balance and density of the SIBA dataset partially compensate for the lower representational capacity of hand-crafted temporal features. Nonetheless, confusion among broadleaved species remains more pronounced than for conifers, consistent with known challenges in deciduous species discrimination.
Overall, the SIBA confusion matrices demonstrate that improved class balance and increased sample size substantially reduce misclassification across all modeling approaches. While fine-tuned deep embeddings provide the most consistent and robust class-level performance, the reduced confusion observed even for simpler models highlights the critical role of data quality in species-level forest classification. These results support the conclusion that targeted data augmentation and balancing strategies can be as influential as model architecture in improving operational NFI mapping performance.
Figure 4: Confusion matrices for Simplified & Imbalanced (SIMB) datasets. Top left: Fine- tune Presto with MLP Classifier, Top right: Presto (Frozen Presto features), Middle left: Fine-tune Presto (Presto features), Middle right: Harmonic + medoid feature, Bottom left: AEF, Bottom right: TESSERA. Each cell has average count and recall percentage.